A flexible pavement is to be designed to last 10 years. The initial PSI is 4.2 and the TSI (the final PSI) is determined to be 2.5. The subgrade has a soil resilient modulus of 15,000 lb/in2 . Reliability is 95% with an overall standard deviation of 0.35. For design, the daily car, pickup truck and light van traffic is 40,000 and the daily truck traffic consists of 1500 passes of single-unit trucks with two single axles and 325 passes of tractor semi-trailer trucks with single, tandem, and triple axles. The axle weights are:

• Cars, pickups, light vans = two 2 kip single axles

• Single-unit truck = 8 kip single axle and 24 kip single axle

• Tractor semi-trailer truck = 10 kip single axle, 18 kip tandem axle, and 42 kip triple axle.
Answer #1

we got D2=3.2 inches. . we will sound it off to hio Ebcz.3.2 would be conservatoire estimate and allow for variations in conssubstituting weget 109.000) =-1.60520-35/19-swi] -020 +10910 [1.772.75 +2.32109 6:5000 -8.07 0.404 1094 (51+195.19 solne for-0.20 Then we have to use the egn. 6 log.Wię = ZRS0 +9:36 (10960 (sN+!) -6.20 +log.[OPSI 29) 0.4D +[1094 (EN +y%) +2.32log meand final Ps] =27 with an initial PSI of 4.2 [ginen in Question]. APSI = 707 TSN = Qi DitQLD, M2 +93 DsM3. 1-0 where 21,92,93for semi tractor tsucks. Note the compositerely small effect of cars and other light vehicles in terms of the 18 kip ESAL Thi[from table] Ciren this, 18 kip equivalent Single arte load for cats, pickups, and light vans is & kip Single axle equivalentSolution * Because the axle - load equivalency factors from standard tables. axle load-equirelency factors are a function of